# VideoGuide: Improving Video Diffusion Models without Training Through a Teacher’s Guide

### v2

@@ 3 VideoGuide
3.1 Video Consistency on Diffusion Trajectory
The DDIM formulation can be expressed as a proximal optimization problem (kim2024dreamsampler):
${\boldsymbol{z}}_{t-1}$ $=\sqrt{\bar{\alpha}_{t-1}}{\boldsymbol{z}}^{\prime}+\sqrt{1-\bar{\alpha}_{t-1}}\hat{\boldsymbol{\epsilon}}_{\theta}({\boldsymbol{z}}_{t},t)\quad\text{where}\quad{\boldsymbol{z}}^{\prime}=\arg\min_{{\boldsymbol{z}}}||{\boldsymbol{z}}-{\boldsymbol{z}}_{0|t}||^{2}_{2}$ (7)
We extend this approach to the video domain by introducing a novel regularization term specially crafted for enhancing temporal consistency.
Specifically, for a given video ${\boldsymbol{x}}_{r}^{1:N}$, suppose that a temporally consistent latent of ${\boldsymbol{z}}_{r}={\mathcal{E}}({\boldsymbol{x}}_{r}^{1:N})$ exists. Then, it would be desirable to set the optimization problem as follows:
$\min_{{\boldsymbol{z}}}||{\boldsymbol{z}}-{\boldsymbol{z}}_{0|t}||^{2}_{2}+\lambda_{reg}R({\boldsymbol{z}})\quad\text{where}\quad R({\boldsymbol{z}}):=||{\boldsymbol{z}}-{\boldsymbol{z}}_{r}||^{2}_{2}$ (8)
Unfortunately, it is infeasible to provide ${\boldsymbol{z}}_{r}$ as the purpose of the VDM is to generate new unseen samples. Thus, we propose to use ${\boldsymbol{z}}_{0|t-\tau}$ as a proxy of ${\boldsymbol{z}}_{r}$ where $\tau$ is a sufficient number of timesteps. This is because ${\boldsymbol{z}}_{0|t-\tau}$ is usually a cleaner and temporally more consistent sample than ${\boldsymbol{z}}_{0|t}$, so we want to utilize this property. Under this assumption, the highly complex problem of generating temporally consistent video samples is reduced to solving the simple optimization problem below:
${\boldsymbol{z}}_{t-1}=\sqrt{\bar{\alpha}_{t-1}}{\boldsymbol{z}}^{\prime}+\sqrt{1-\bar{\alpha}_{t-1}}\hat{\boldsymbol{\epsilon}}_{\theta}({\boldsymbol{z}}_{t},t)$ (9)
$\text{where}\quad{\boldsymbol{z}}^{\prime}=\min_{{\boldsymbol{z}}}||{\boldsymbol{z}}-{\boldsymbol{z}}_{0|t}||^{2}_{2}+\lambda_{reg}||{\boldsymbol{z}}-{\boldsymbol{z}}_{0|t-\tau}||^{2}_{2}$
which is equivalent to
${\boldsymbol{z}}_{t-1}=\sqrt{\bar{\alpha}_{t-1}}\left(\beta{\boldsymbol{z}}_{0|t}+(1-\beta){\boldsymbol{z}}_{0|t-\tau}\right)+\sqrt{1-\bar{\alpha}_{t-1}}\hat{\boldsymbol{\epsilon}}_{\theta}({\boldsymbol{z}}_{t},t)$ $,\quad\beta=\frac{1}{1+\lambda_{reg}}$ (10)
Accordingly, it suffices to use the interpolation of ${\boldsymbol{z}}_{0|t}$ and ${\boldsymbol{z}}_{0|t-\tau}$ as an estimate of the temporally consistent form of ${\boldsymbol{z}}_{0|t}$. To further ensure model flexibility to refine high-frequency areas for better image quality, we employ a low-pass filter inspired by previous work (wu2023freeinit). Specifically, using a low-pass filter and high-pass filter of cut-off frequency $\gamma$, denoted $\textit{LPF}_{\gamma}$ and $\textit{LPF}_{\gamma}$ respectively, we define the following update:
${\boldsymbol{z}}_{t-1}=\textit{LPF}_{\gamma}({\boldsymbol{z}}_{t-1})+\textit{HPF}_{1-\gamma}({\boldsymbol{\epsilon}})\quad\text{where}\quad{\boldsymbol{\epsilon}}\sim N(0,{\mathbf{I}})$ (11)
Replacement of high-frequency regions with random Gaussian noise enhances model capacity to infer corresponding high-frequency components, leading to denoised results of higher quality.
3.2 Guidance with External Video Diffusion Models
@@
The assumption ${\boldsymbol{z}}_{r}\approx{\boldsymbol{z}}_{0|t-\tau}$ in Sec. 3.1 holds for any sample ${\boldsymbol{z}}_{0|t-\tau}$ with temporal consistency comparable to a real-world sample. This brings us to realize that the sample ${\boldsymbol{z}}_{0|t-\tau}$ does not necessarily have to originate from the same base model. It is possible to plug in any denoised latent ${\boldsymbol{z}}_{0|t-\tau}$ from any video diffusion model, and the denoising process would be guided to follow the temporal consistency of the supplemented latent. Here, we demonstrate the steps required for utilizing denoised samples ${\boldsymbol{z}}_{0|t-\tau}^{(G)}$ of an external guidance model $G$ to enhance the performance of the base sampling model $S$.
@@
Renoising into the Guidance Domain. Different video diffusion models are trained on different noise schedules and distributions, and matching such discrepancies is a mandatory process. When utilizing a guiding model with conflicting factors, the intermediate latent ${\boldsymbol{z}}_{t}$ of the sampling model must be transformed to align with the noise schedule and distribution of the guiding model. The transformation process can be defined as follows:
${\boldsymbol{z}}_{t}^{(G)}=\sqrt{\bar{\alpha}_{t}^{(G)}}{\boldsymbol{z}}^{(S)}_{0|t}+\sqrt{1-\bar{\alpha}^{(G)}_{t}}{\boldsymbol{\epsilon}},\quad\mbox{where}\quad{\boldsymbol{\epsilon}}\sim N(0,{\mathbf{I}})$ (12)
@@
where $(S)$ denotes the components related to the base sampling model and $(G)$ denotes the components related to the external guiding model. Specifically, ${\boldsymbol{z}}^{(S)}_{0|t}$ is the denoised sample from ${\boldsymbol{z}}^{(S)}_{t}$ at timestep $t$, and $\bar{\alpha}_{t}^{(G)}$ is derived from the noise schedule of the guiding diffusion model. The resulting outcome ${\boldsymbol{z}}_{t}^{(G)}$ can then be denoised with the guiding model for a sufficient number of timesteps $\tau$ up to ${\boldsymbol{z}}_{0|t-\tau}^{(G)}$.
@@
${\boldsymbol{z}}_{t-1}^{(S)}=\sqrt{\bar{\alpha}_{t-1}}(\beta{\boldsymbol{z}}_{0|t}^{(S)}+(1-\beta){\boldsymbol{z}}_{0|t-\tau}^{(G)})+\sqrt{1-\bar{\alpha}_{t-1}}\hat{\boldsymbol{\epsilon}}_{\theta}^{(S)}({\boldsymbol{z}}_{t},t)$ (13)
@@
Note that the only difference from Eq. (10) is the introduction of the ${\boldsymbol{z}}_{0|t-\tau}^{(G)}$ term, where originally ${\boldsymbol{z}}_{0|t-\tau}^{(S)}$ would be used. $\textit{LPF}_{\gamma}$ can then be used on $z_{t-1}^{(S)}$ as in Eq. (11) for replacing high-frequency components:
${\boldsymbol{z}}_{t-1}^{(S)}=\textit{LPF}_{\gamma}({\boldsymbol{z}}^{(S)}_{t-1})+\textit{HPF}_{1-\gamma}({\boldsymbol{\epsilon}})\quad\text{where}\quad{\boldsymbol{\epsilon}}\sim N(0,{\mathbf{I}})$ (14)
Synergistic Effects of External VDM Guidance. Utilizing a high-performance open-source model (chen2024videocrafter2) as the guiding diffusion model in our VideoGuide framework is shown to improve temporal consistency even while achieving faster convergence. Compared to the self-guided case, generating temporally coherent samples from a superior model proves beneficial to the quality of the resulting samples, as illustrated in Sec. 4. Moreover, since interpolation occurs only during the early timesteps, the advantages of the sampling diffusion model—such as the personalized video generation and controllability of AnimateDiff—are fully preserved. Accordingly, VideoGuide is a versatile framework that can combine the best of both worlds: the sampling model and the guiding model. No additional training or fine-tuning is required for seeing such synergistic effects, allowing the user to freely select favored VDMs in a plug-and-play fashion.
3.3 VideoGuide in Practice
Early Timestep Interpolation. In VideoGuide a novel interpolation technique is included in the inference process, and the equations above explain cases at a specific timestep $t$. Theoretically this interpolation could be performed at every denoising timestep, but such iteration would both be computationally expensive and detrimental to the high-frequency components that emerge at later timesteps. Recent work (wu2023freeinit) shows that providing informative low-frequency components at initialization time is sufficient for enhancing temporal consistency. Likewise, we find that applying our interpolation scheme at early timesteps is adequate for enforcing temporal consistency while allowing high-frequency regions to align more closely to the low-frequency structure. An extensive ablation study regarding the number of interpolation steps is given in Sec. 5.1
@@ 3.3 VideoGuide in Practice
Prior Distillation. Each video diffusion model spans its own specific data distribution, causing sample generation to be restricted to the data prior the model has been trained on. Thus, if the data prior of a model is substandard, the generation results of the model are also inherently substandard. This is especially noticeable when using personalized text-to-image (T2I) models such as Dreambooth or LoRA in AnimateDiff, in which substandard results that do not align with the given text prompt are frequently observed. Prior work (ge2024preservecorrelationnoiseprior) elaborates on the importance of data prior for VDMs, but the proposed solution involves extensive fine-tuning, making it impractical for simple use cases. On the other hand, VideoGuide comes as a potential solution in such cases, where the interpolation between two models exhibit a form of prior distillation. Through the guidance of a generalized video diffusion model (e.g. chen2024videocrafter2) the base sampling model is able to refer to the denoised sample provided by the guidance model, and steer its sampling process towards a relevant outcome. This allows for the effective generation of diverse objects, even while retaining the style of the original data domain. For the case of AnimateDiff, the approach allows for broader customization without the need for directly training the personalized T2I model on a wider range of data. Extensive analysis concerning this issue is provided in Sec. 5.2.
@@ 4 Experiments
Experimental Settings. In our experiments, we leverage multiple open-source Text-to-Video (T2V) diffusion models to explore the combined strengths of each. For the guiding diffusion model, we choose Videocrafter2 (chen2024videocrafter2) due to its strong performance in temporal consistency, as measured by the VBench (huang2023vbench) benchmark. For sampling, we employ AnimateDiff (guo2023animatediff) for flexible personalization of video content, and Lavie (wang2023lavie) to enhance video quality and increase frame count through super-resolution and interpolation techniques. This integration combines the temporal consistency of the guiding model with the advantages of the sampling model. All experiments were conducted using DDIM with 50 steps for sampling. For our experiments with AnimateDiff, we set $I=5$, $\beta=0.5$, and $\tau=10$, and used the Butterworth filter with a normalized frequency of $0.25$ and a filter order of $n=4$. Additional experimental details are provided in Appendix A.
Evaluation Metrics. To validate the improvement in video consistency with our proposed method, we evaluate four key metrics: subject consistency, background consistency, imaging quality, and motion smoothness. For subject consistency evaluation, DINO (caron2021emerging) feature similarity between frames is measured to assess consistency of the subject’s appearance throughout the video. Background consistency is evaluated using CLIP feature similarity between frames to evaluate overall scene consistency. Imaging quality is also a key metric in that maintaining original image quality is essential for generation and enabling customization. Thus we evaluate image quality using the multi-scale image quality transformer (MUSIQ) (ke2021musiq), which measures frame-wise low-level distortion such as noise, blur, and over-exposure. Additionally, to ensure smooth motion, we employ a video interpolation model (licvpr23amt) to assess consistency of motion across video frames.
Figure 3: Qualitative Comparison. VideoGuide is applied on various base models for different text prompts. For each prompt, frames of generated samples from four different models are displayed: (i) Base model (first row); (ii) Base model with FreeInit (second row); (iii) Base model with VideoGuide (self-guided case) (third row); (iv) Base model with VideoGuide (external model-guided case) (fourth row). AD, VC, LV indicate guidance models of AnimateDiff, VideoCrafter-2.0, LaVie, respectively. Samples for the base model show substandard temporal consistency, especially regarding color fluctuation and subject appearance change. Applying FreeInit improves consistency but introduces degradation in imaging quality, such as smoothing out of textural details. In contrast, applying VideoGuide significantly enhances temporal consistency while preserving imaging quality, both for the self-guided case and the external model-guided case.

### v3

@@ 3 VideoGuide
In this section, we present VideoGuide, a novel guiding framework which improves the video consistency without compromising significant computational costs in contrast to prior works. Our framework is based on a teacher-guided latent optimization objective, which when minimized during the early stage of reverse sampling process, progressively improves the temporal consistency of generated video. We begin by outlining the overall optimization framework.
3.1 Video Consistency Guidance
An important contribution of this work is revealing that video consistency can be enhanced by recasting guidance as an optimization problem. We introduce a new objective that regularizes the sampling path of the reverse diffusion process to improve the quality of generated video samples:
$\ell({\bm{z}}_{0};\psi,{\bm{\epsilon}},t)=||{\bm{\epsilon}}_{\psi}(\sqrt{\bar{\alpha}_{t}}{\bm{z}}_{0}+\sqrt{1-\bar{\alpha}_{t}}{\bm{\epsilon}},t)-{\bm{\epsilon}}||^{2}_{2},$ (7)
where $\psi$ conceptually represents a general teacher model which can denoise the noisy video latents reasonably well. That said, this regularizer represents an ideal condition that high-quality video samples should satisfy: the ideal in-distribution video samples should be well reconstructed from random perturbations followed by denoising using teacher video diffusion models.
Then, we can now integrate the optimization step of (7) in reverse sampling process as a guidance in terms of denoised video estimates ${\bm{z}}_{0|t}$. This reads a single DDIM iterate as follows:
${\bm{z}}_{t-1}$ $=\sqrt{\bar{\alpha}_{t-1}}({\bm{z}}_{0|t}-\gamma_{t}\nabla_{z_{0|t}}\ell({\bm{z}}_{0|t};\psi,{\bm{\epsilon}},t))$ (8)
$\quad+\sqrt{1-\bar{\alpha}_{t-1}}{\bm{\epsilon}}_{\theta}({\bm{z}}_{t},t)$
Observe that Eq. (7) can be reformulated by swapping ${\bm{\epsilon}}_{\psi}({\bm{z}}_{t},t)$ into an expression of ${\bm{z}}_{t}$ and ${\bm{z}}_{0|t}^{\psi}$, and swapping ${\bm{\epsilon}}$ into an expression of ${\bm{z}}_{t}$ and ${\bm{z}}_{0}$ as follows:
$\ell({\bm{z}}_{0};\psi,{\bm{\epsilon}},t)$ $=\left\|\frac{{\bm{z}}_{t}-\sqrt{\bar{\alpha}_{t}}{\bm{z}}_{0|t}^{\psi}}{\sqrt{1-\bar{\alpha}_{t}}}-\frac{{\bm{z}}_{t}-\sqrt{\bar{\alpha}_{t}}{\bm{z}}_{0}}{\sqrt{1-\bar{\alpha}_{t}}}\right\|^{2}_{2}$ (9)
$=\frac{\bar{\alpha}_{t}}{1-\bar{\alpha_{t}}}\left\|{\bm{z}}_{0}-{\bm{z}}_{0|t}^{\psi}\right\|^{2}_{2}$
Plugging Eq. (9) into Eq. (8), the gradient term dissolves into an interpolation scheme and the update rule can be simplified as follows:
${\bm{z}}_{t-1}=\sqrt{\bar{\alpha}_{t-1}}{\bm{z}}^{\prime}+\sqrt{1-\bar{\alpha}_{t-1}}{\bm{\epsilon}}_{\theta}({\bm{z}}_{t},t)$ (10)
$\text{where}\quad{\bm{z}}^{\prime}=\beta\cdot{\bm{z}}_{0|t}+(1-\beta)\cdot{\bm{z}}_{0|t}^{\psi}$
with $\beta=1-\frac{2\gamma_{t}\bar{\alpha}_{t}}{1-\bar{\alpha}_{t}}$. ${\bm{z}}_{0|t}^{\psi}$ can be obtained with the renoising process of ${\bm{z}}_{0|t}$ as below:
${\bm{z}}_{0|t}^{\psi}=\frac{{\bm{z}}_{t}-\sqrt{1-\bar{\alpha}_{t}}{\bm{\epsilon}}_{\psi}({\bm{z}}_{t},t)}{\sqrt{\bar{\alpha}_{t}}}$ (11)
where ${\bm{z}}_{t}=\sqrt{\bar{\alpha}_{t}}{\bm{z}}_{0|t}+\sqrt{1-\bar{\alpha_{t}}}{\bm{\epsilon}},\quad{\bm{\epsilon}}\sim N(0,{\mathbf{I}})$
Specifically, a key challenge in our approach is the lack of a direct model for $\psi$, which, ideally, would take the form of a consistency model [song2023consistency]. To address this, we approximate $\psi$ using an iterative reverse sampling method to predict the endpoint of the probability flow ODE (PF-ODE). By performing reverse sampling over multiple steps $\tau$ in the original model $\theta$, we generate samples that serve as a proxy for the PF-ODE endpoint. Consequently, we can redefine the terms as follows:
${\bm{z}}_{0|t}^{\psi}\approx{\bm{z}}_{0|t-\tau},\quad{\bm{z}}^{\prime}=\beta\cdot{\bm{z}}_{0|t}+(1-\beta)\cdot{\bm{z}}_{0|t-\tau}$ (12)
This approach offers a cost-effective solution for approximating $\psi$ in video diffusion models, eliminating the need to train a separate consistency model.
To address the increased sampling time caused by interpolating multiple $\tau$ samples at each reverse step, we propose applying a low-frequency filter during the early timesteps of the diffusion process. Recent work [yu2024scaling] indicates that these initial stages primarily establish low-frequency structures, with high-frequency components contributing minimally to image quality. By introducing a low-frequency filter early in the diffusion trajectory, we can accelerate convergence toward consistent samples without sacrificing quality.
Unlike previous methods [wu2023freeinit], which apply the filter only to the initial noise, our approach iteratively applies the filter along the entire diffusion path, ensuring stability throughout the trajectory. Additionally, by incorporating early stopping and leveraging our filter’s ability to rapidly achieve sample consistency, we prevent the image degradation typically observed with prolonged optimization [kim2023collaborative, hertz2023deltadenoisingscore] of score distillation sampling (SDS) loss [poole2022dreamfusion], which is similar to our objective. Specifically, we define our updates using low-pass ($\textit{LPF}_{\gamma}$) and high-pass ($\textit{HPF}_{\gamma}$) filters with a cutoff frequency $\gamma$, streamlining the sampling process while maintaining high quality.
${\bm{z}}_{t-1}=\textit{LPF}_{\gamma}({\bm{z}}_{t-1})+\textit{HPF}_{1-\gamma}({\bm{\epsilon}})$ (13)
$\text{where}\quad{\bm{\epsilon}}\sim N(0,{\mathbf{I}})$
3.2 Guidance with External VDMs
@@
The assumption of $\psi$ in Sec. 3.1 holds for any teacher model that provides a reliable estimate of the denoised sample. This brings us to realize that $\psi$ does not necessarily have to be approximated by the same base model. It is possible to plug in any video diffusion model to approximate $\psi$ and the denoising process would be guided to follow the temporal consistency of the supplemented latent. Here, we demonstrate the steps required for utilizing denoised samples ${\bm{z}}_{0|t-\tau}^{(G)}$ of an external guidance model $G$ to enhance the performance of the base sampling model $S$.
@@
Renoising into the Guidance Domain. Different video diffusion models are trained on varying noise schedules and distributions, and matching such discrepancies is a mandatory process. When utilizing a guiding model with conflicting factors, the intermediate latent ${\bm{z}}_{t}$ of the sampling model must be transformed to align with the noise schedule and distribution of the guiding model. The transformation process can be defined as follows:
${\bm{z}}_{t}^{(G)}=\sqrt{\bar{\alpha}_{t}^{(G)}}{\bm{z}}^{(S)}_{0|t}+\sqrt{1-\bar{\alpha}^{(G)}_{t}}{\bm{\epsilon}}$ (14)
$\text{where}\quad{\bm{\epsilon}}\sim N(0,{\mathbf{I}})$
@@
where $(S)$ denotes the components related to the base sampling model and $(G)$ denotes the components related to the external guiding model. Specifically, ${\bm{z}}^{(S)}_{0|t}$ is the denoised sample from ${\bm{z}}^{(S)}_{t}$ at timestep $t$, and $\bar{\alpha}_{t}^{(G)}$ is derived from the noise schedule of the guiding diffusion model. The resulting outcome ${\bm{z}}_{t}^{(G)}$ can then be denoised with the guiding model for a sufficient number of timesteps $\tau$ up to ${\bm{z}}_{0|t-\tau}^{(G)}$. Thus, the PF-ODE endpoint originally approximated by $\psi$ is now represented as ${\bm{z}}_{0|t-\tau}^{(G)}$, the denoised output of the guiding model $G$.
@@
${\bm{z}}_{t-1}^{(S)}$ $=\sqrt{\bar{\alpha}_{t-1}}(\beta\cdot{\bm{z}}_{0|t}^{(S)}+(1-\beta)\cdot{\bm{z}}_{0|t-\tau}^{(G)})$ (15)
$\quad+\sqrt{1-\bar{\alpha}_{t-1}}{\bm{\epsilon}}_{\theta}^{(S)}({\bm{z}}_{t},t)$
@@
Note that the only difference from Eq. (10) is the introduction of the ${\bm{z}}_{0|t-\tau}^{(G)}$ term, where originally ${\bm{z}}_{0|t-\tau}^{(S)}$ would be used. $\textit{LPF}_{\gamma}$ can then be used on ${\bm{z}}_{t-1}^{(S)}$ as in Eq. (13) for replacing high-frequency components:
${\bm{z}}_{t-1}^{(S)}=\textit{LPF}_{\gamma}({\bm{z}}^{(S)}_{t-1})+\textit{HPF}_{1-\gamma}({\bm{\epsilon}})$ (16)
$\text{where}\quad{\bm{\epsilon}}\sim N(0,{\mathbf{I}})$
In cases where external video diffusion models (VDMs) are used as guidance, the low-frequency filter serves an additional role in preserving domain fidelity. By controlling high-frequency components and introducing Gaussian noise, it prevents unwanted domain drift toward the guidance model while selectively distilling only its temporal stability. This approach captures the temporal consistency of the guiding diffusion model while preserving unique characteristics, such as those in AnimateDiff, without compromising image quality. Notably, these synergistic effects are achieved without additional training or fine-tuning, enabling users to flexibly employ preferred video diffusion models (VDMs) in a plug-and-play manner.
3.3 Prior Distillation
@@
Each video diffusion model spans its own specific data distribution, causing sample generation to be restricted to the data prior the model has been trained on. Thus, if the data prior of a model is substandard, the generation results of the model are also inherently substandard. This is especially noticeable when using personalized text-to-image (T2I) models such as Dreambooth or LoRA in AnimateDiff, in which substandard results that do not align with the given text prompt are frequently observed. Prior work [ge2024preservecorrelationnoiseprior] elaborates on the importance of data prior for VDMs, but the proposed solution involves extensive fine-tuning, making it impractical for simple use cases. On the other hand, VideoGuide comes as a potential solution in such cases, where the interpolation between two models exhibit a form of prior distillation. Through the guidance of a generalized video diffusion model (e.g. [chen2024videocrafter2]) the base sampling model is able to refer to the denoised sample provided by the guidance model, and steer its sampling process towards a relevant outcome. This allows for the effective generation of diverse objects, even while retaining the style of the original data domain. For the case of AnimateDiff, the approach allows for broader customization without the need for directly training the personalized T2I model on a wider range of data. Extensive analysis concerning this issue is provided in Sec. 5.2.
@@ 4 Experiments
Experimental Settings. In our experiments, we leverage multiple open-source Text-to-Video (T2V) diffusion models to explore the combined strengths of each. For the guiding diffusion model, we choose Videocrafter2 [chen2024videocrafter2] due to its strong performance in temporal consistency, as measured by the VBench [huang2023vbench] benchmark. For sampling, we employ AnimateDiff [guo2023animatediff] for flexible personalization of video content, and Lavie [wang2023lavie] to enhance video quality and increase frame count through super-resolution and interpolation techniques. This integration combines the temporal consistency of the guiding model with the advantages of the sampling model. All experiments were conducted using DDIM with 50 steps for sampling. For our experiments with AnimateDiff, we set the number of interpolation steps $I=5$, $\beta=0.5$, and $\tau=10$, and used the Butterworth filter with a normalized frequency of $0.25$ and a filter order of $n=4$. Additional experimental details are provided in Supplementary Material.
Evaluation Metrics. To validate the improvement in video consistency with our proposed method, we evaluate five key metrics: subject consistency, background consistency, imaging quality, motion smoothness, and dynamic degree. For subject consistency evaluation, DINO [caron2021emerging] feature similarity between frames is measured to assess consistency of the subject’s appearance throughout the video. Background consistency is evaluated using CLIP feature similarity between frames to evaluate overall scene consistency. Imaging quality is also a key metric in that maintaining original image quality is essential for generation and enabling customization. Thus we evaluate image quality using the multi-scale image quality transformer (MUSIQ) [ke2021musiq], which measures frame-wise low-level distortion such as noise, blur, and over-exposure. To ensure smooth motion, we employ a video interpolation model [licvpr23amt] to assess consistency of motion across video frames. To compare the magnitude of the motion in the videos, we utilize RAFT [teed2020raftrecurrentallpairsfield] to estimate the optical flow, and calculate the mean of top 5% largest magnitude of the flows.